OpenStudy (anonymous):

I really need help. Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem.

OpenStudy (anonymous):

x(.50 + 40%) = 500

OpenStudy (anonymous):

or x(0.50 + 0.40) = 500 if your teacher wants both to be in decimals.

OpenStudy (anonymous):

so how many flowers would they need?

OpenStudy (anonymous):

Solve the equation. 0.50 plus 0.40 = 0.70. Your new equation is 0.70x = 500 Solve for x by dividing 500 by 0.70

OpenStudy (anonymous):

You should get an answer around 714

OpenStudy (anonymous):

thats what I got . But everyone else is getting the answer of 2,500 .

OpenStudy (anonymous):

Ummm. They might be right. You will need 714 flowers to get 500 dollars TO GIVE to the foundation. With this you aren't getting any profit.

OpenStudy (anonymous):

But i don't understand how their answer is to high. I wouldve thought to just mulitply the number of flowers by two and get 1,400 not 2,500

OpenStudy (anonymous):

this is how they got it.. i just dont understand the way they put it.. P = the profit they make n = the number of flowers they sell m = the mark-up of the flowers c = the cost of a flower Then P = mcn - cn Rearrange to find n P = n(mc - c) n = P / (mc - c) Substituting values n = 500 / (1.4*0.5 - 0.5) n = 2500 They must sell 2500 flowers to make a profit of $500.

OpenStudy (anonymous):

I see what they did. They arranged the problem in order to find how many flowers are needed so that the 40% adds up to 500 dollars.

OpenStudy (anonymous):

So how would you write the equation?

OpenStudy (anonymous):

The equation can be written in two ways: 1. P = mcn - cn or 2. 500 = (40%)(.50)n - .50n

OpenStudy (anonymous):

Okay, thank you!(:

OpenStudy (anonymous):

haha no problem love c: